OpenStudy (anonymous):

Simplify the following expression. 2a+3-9b-a-7

OpenStudy (anonymous):

(2a-a)-9b+(3-7)

OpenStudy (anonymous):

combine like terms... 2a - a = a -9b is all by itself -7 + 3 = - 4 so now you have : a - 9b - 4